### v1

# Human Mesh Modeling for Anny Body

Abstract
Parametric body models are central to many human-centric tasks, yet existing models often rely on costly 3D scans and learned shape spaces that are proprietary and demographically narrow. We introduce Anny, a simple, fully differentiable, and scan-free human body model grounded in anthropometric knowledge from the MakeHuman community. Anny defines a continuous, interpretable shape space, where phenotype parameters (e.g. gender, age, height, weight) control blendshapes spanning a wide range of human forms — across ages (from infants to elders), body types, and proportions. Calibrated using WHO population statistics, it provides realistic and demographically grounded human shape variation within a single unified model. Thanks to its openness and semantic control, Anny serves as a versatile foundation for 3D human modeling — supporting millimeter-accurate scan fitting, controlled synthetic data generation, and Human Mesh Recovery (HMR). We further introduce Anny-One, a collection of 800k photorealistic humans generated with Anny, showing that despite its simplicity, HMR models trained with Anny can match the performance of those trained with scan-based body models, while remaining interpretable and broadly representative. The Anny body model and its code are released under the Apache 2.0 license, making Anny an accessible foundation for human-centric 3D modeling.
1 Introduction
Understanding human shape and motion in 3D is fundamental to computer vision, graphics, and robotics. Parametric body models provide a compact and differentiable representation of the human body, enabling a wide range of applications — from scan fitting and animation to Human Mesh Recovery (HMR) from images and videos [goel2023humans, yin2025smplest, zhang2021pymaf, shin2024wham, kocabas2020vibe]. These models have been central to progress in human-centric perception, powering tasks such as motion analysis, behavior understanding, and human–robot interaction.
Among them, the SMPL family [loper2015smpl, pavlakos2019smplx, osman2020star, osman2022supr] has played a transformative role. SMPL-like models achieve high accuracy on standard benchmarks by learning low-dimensional shape spaces from 3D body scans [robinette2002caesar, Bogo:CVPR:2014, Bulat2017FAW]. However, this data-driven design also imposes limitations: the available scan datasets are expensive, demographically narrow, and privacy-sensitive. As a result, current models struggle to represent the full diversity of human bodies—especially children, elderly individuals, and morphologies uncommon in populations. Separate models have been proposed for specific demographics, such as SMIL [hesse2018smil] for infants or distinct male/female SMPL variants, and more recent efforts like SMPL-A [agora] interpolate between children and adults. Yet, obtaining representative data for all body types remains impractical and privacy-sensitive.
Figure 1: Shape parametrization is implemented using piece-wise multi-linear interpolation between prototypical shapes (illustration with age).
Instead of relying on scan data, we explore an alternative foundation for human modeling. We leverage the anthropometric knowledge embedded in MakeHuman [makehuman]—a free, community-driven framework designed to model human variability through explicit, interpretable parameters such as age, gender, and body proportions. This procedural knowledge, built by artists over decades, provides a rich, open description of human morphology that naturally spans diverse body types.
Building on this foundation, we introduce Anny, a simple and fully differentiable human body model that replaces learned shape bases with interpretable phenotype parameters. Each parameter (e.g., age, height, weight, muscle, proportions) is defined in the continuous range [0,1], and directly controls corresponding blendshapes (Figure 1). Calibrated using WHO population statistics, Anny covers realistic human variation across the full lifespan—from infants to elders—within a single unified model (Figure Human Mesh Modeling for Anny Body). Because it is built on open assets rather than biometric scans, Anny is free from privacy constraints and can be shared, analyzed, and extended by the community.
Refer to captionRefer to captionRefer to caption
head-fat
Refer to captionRefer to captionRefer to caption
upperleg-height
Refer to captionRefer to caption
stomach-pregnant
Figure 2: Example of local morphological changes covered by the model.
Because it is open and semantically defined, Anny can be easily inspected, extended, and integrated into existing pipelines. This design enables a wide range of applications — from millimeter-accurate scan fitting to controllable synthetic data generation and HMR — within a single unified framework. To demonstrate its potential, we create Anny-One, a large-scale synthetic dataset of 800k photorealistic humans generated with Anny, featuring expressive full-body poses, hands, and faces across diverse environments. HMR models trained with Anny-One achieve competitive accuracy on standard benchmarks and outperform existing approaches when body-shape diversity is high — using a single parametric model.
Our main contributions are: (A) Anny, a differentiable, scan-free human body model representing continuous and interpretable shape variations across genders, ages, and body types. (B) Anny-One, a large-scale synthetic dataset of 800k photorealistic humans with diverse 3D poses and shapes; (C) Empirical validation showing that Anny enables accurate scan fitting and competitive HMR, offering a unified and complete approach to human modeling. This work advocates for moving beyond scan-driven latent spaces toward open, semantically meaningful, and globally representative human models. Code will be released under the Apache 2.0 license to encourage accessible foundation tool for human-centric 3D modeling.
2 Related work
Modeling people using 3D human meshes is an active area of research in computer vision, motivated by the fact that the way humans behave and interact with their environment is better captured when modeling the full 3D surface of the body. Differentiable body models, in particular, can be integrated with vision backbones, serving as a structured output representation that enables efficient HMR from input images or videos.
Surface mesh models
have been used to model the appearance of objects since the early days of computer graphics. Skeleton-based deformations and shape blending have been a standard way to model variability in pose and shape for decades [lewis2000posespace], and are particularly suited for human characters. Important research efforts have been made to build accurate human models representative of body surface deformations across different poses and across different individuals characteristics, through data-driven approaches based on 3D human scans. The SMPL [loper2015smpl] body model family is arguably the most widely used in the computer vision community. It relies on linear blend shapes to model shape-dependent vertex displacements in rest pose, as well as pose-corrective blend shapes to further refine the output mesh depending on the pose. The original model was extended to SMPL-X [pavlakos2019smplx] to capture full-body meshes, i.e., including facial expressions and hand poses. Other refinements were also considered, in particular STAR [osman2020star] proposed a more compact formulation compared to SMPL, and SUPR [osman2022supr] split the model into individual body parts with a sparse factorization of pose-corrective blend shapes. To model infants, a separate model SMIL [hesse2018smil] was also proposed, sometimes combined with SMPL through interpolation, referred-to as SMIL-X [agora]. Other body models with similar properties have been proposed, such as GHUM and GHUML [xu2020ghum] that use the latent space of an auto-encoder as parametrization space. More recently, ATLAS [park2025atlas] introduced a high-fidelity, scan-based model that decouples surface and skeletal representations for finer control over body proportions.
Shape diversity.
A major challenge for parametric body models is to capture the diversity of body shapes observed around the world. This is hard to achieve because collecting data such as 3D body scans is costly and time consuming. Yet the above approaches, namely the SMPL and GHUM families, tackle this by learning shape representations from datasets of 3D scans. In particular, the CAESAR [robinette2002caesar] dataset used to train SMPL comprises people from the USA, Netherlands and Italy aged between 18 and 65 years in early 2000s. It contains fewer than $5000$ individuals, and already constitutes a significant data collection effort. This introduces limits on body models fitted from data. We instead rely on expert artistic manual definitions of the human body. Accurately modeling the distribution of human shape has been the focus of much efforts from computer graphics designers and artists; we leverage these efforts instead of collecting real-world data.
Various tools exist to design human-like characters [makehuman, humgen3d, metahuman]. We build our work on MakeHuman [makehuman], an open-source and community-driven project. In combination with a synthetic data generation pipeline, we obtain an efficient way of generating diverse body shapes at scale. Empirically, we find that this is sufficient to obtain competitive HMR results, while avoiding costly 3D body scans.
Human mesh recovery,
as introduced in the eponymous work HMR [kanazawa2018end], aims to estimate a human mesh given an image. Here, we focus on parametric HMR, which refers to models that output the parameters of a parametric human body model such as SMPL [loper2015smpl]. Existing methods can be broadly separated in two categories: single-person and multi-person methods. Single-person methods assume that detections or bounding boxes for humans are provided, either from ground truth data or by an off-the-shelf detection model. In that case the HMR model directly regresses body model parameters [kanazawa2018end]. In that setting, progress was made at the level of architectures and backbones [goel2023humans, armando2024cross], or camera models [wang2023zolly, patel2025camerahmr, wang2024blade], and extensions to expressive body models [multi-hmr2024, cai2023smpler, lin2023one, sun2024aios, feng2021collaborative]. In multi-person HMR methods [multi-hmr2024, sun2021monocular, sun2024aios, sun2022putting], all humans in the input have to be detected and placed in the 3D scene by the model, in addition to regressing each human mesh. In that case, the input domain of the model typically shifts from center crops of people to images containing multiple people at arbitrary locations. Our proposed Anny-One dataset is suitable for both settings, and we empirically evaluate it with both single-person and multi-person state-of-the-art methods.
Training data for HMR.
Another avenue of research to improve the performance of HMR models is to focus on obtaining better training data. One of the main limiting factors in training HMR models is the lack of images labeled with 3D ground-truth humans. While earlier works [pavlakos2019smplx, bogo2016keep, huang2017towards, lassner2017unite, guler2018densepose] used optimization procedures to fit parametric models to 2D observations such as 2D keypoints, most regression-based models cited above use datasets with 2D [lin2014microsoft, andriluka14cvpr, johnson2010clustered] or 3D ground-truth joints [ionescu2013human3, mono-3dhp2017], sometimes with pseudo-ground-truth meshes [li2022cliff, joo2021exemplar, kolotouros2019learning, moon2022neuralannot, lin2023one]. Acquiring images with 3D ground truth is expensive, time-consuming, and pseudo-annotations often lack precision in particular for hands and faces. A popular solution is to use synthetic data; it presents 3 major advantages: 1) scaling is efficient 2) annotations are free of noise 3) the diversity of the data, in terms of many variables such as body shapes, skin tones, ages, motions, environments, or cameras is easy to control. Existing works have shown that synthetic data can complement real data [varol2017learning, agora], and more recently that synthetic data alone leads to state-of-the-art (SOTA) HMR models [bedlam, yin2024whac, multi-hmr2024], in particular when expressive poses (including hands and faces) are involved [multi-hmr2024] or accurate camera estimation in human-centric scenes is required [wang2024blade, patel2025camerahmr].
In this work, we leverage all advantages of synthetic data. Our large-scale Anny-One dataset contains diverse humans in terms of body shape and appearance, with expressive poses including faces and hands, with exact ground truth. We train expressive single-person and multi-person HMR models, and obtain SOTA performance without using real images for training.
3 Modeling Anny Body
We propose Anny, a differentiable parametric mesh model aiming at modeling a large diversity of human morphologies. Anny is built upon assets from MakeHuman [makehuman, mpfb2], a community-driven framework available under an open-source license that enables artists to model a wide variety of human-like 3D characters. We release a PyTorch implementation of Anny under a permissive Apache 2.0 open-source license to foster the development of human-centric research and applications.
Base model.
The base model of Anny is composed of 13,380 vertices and 13,378 quadrilateral faces (excluding tongue and eyes), assigned with weights to a default rig of 163 bones, illustrated Fig. 3. Other rigs are additionally supported, such as the one typically used in Mixamo [mixamo] character animations, which allows for simple animation retargeting by reproducing bone orientations.
Figure 3: Skeleton of Anny. Left: default skeleton shown for different morphologies, excluding facial bones for visualization. Right: use of the same skeleton as a Mixamo [mixamo] character animation for retargeting purposes.
Phenotypes.
Following MakeHuman terminology, we capture the diversity of human morphology through various parameters, which we refer to as phenotypes. These phenotypes aim to encode high-level characteristics, such as age (Fig. 1), gender, weight, muscle amount, etc., as well as more local changes, such as the amount of head fat or changes in belly morphology during pregnancy (Fig. 2). Phenotypes are grounded into a set of prototypical blendshapes, each modeling a 3D human mesh with particular characteristics, e.g. a female baby having small muscle, average weight, and large height.
Word of caution.
Phenotypes are based on preconceptions of artists regarding particular human traits. As a result, they encode by design stereotypes of MakeHuman artists, and one should not expect phenotype parameters to faithfully encode any identity-related characteristics, such as gender, age or ethnicity. We nonetheless found these parameters useful to model the diversity of human morphologies, and we chose to keep MakeHuman terminology to make explicit the intent of the artists. Existing data-driven models [loper2015smpl, pavlakos2019smplx, osman2020star, xu2020ghum, osman2022supr, park2025atlas] avoid addressing this delicate question by using abstract shape spaces with no explicit semantics–except for gender-specific models. They nonetheless also convey biases related to their topology (all mesh models, including Anny, assume individuals with four limbs) and to their training distribution (most models are only suited for adults, for example).
Figure 4: Model calibration. Calibration of Anny morphological shape distribution with WHO Child Growth standards for boys [who2006growth].
Differentiable deformation.
Anny takes shape and pose parameters as input and outputs a 3D posed human mesh representation in a backward-differentiable manner. The shape of the mesh is controlled by scalar coefficients targeting various phenotypes regarding age, weight, gender, etc.. These coefficients define weights of a piecewise-multilinear interpolation between prototypical blendshapes, that are used to adjust the shape of a base model. For instance, blendshapes corresponding to the phenotypes child (age=1/3age{=}1/3) and young (age=2/3age{=}2/3) contribute equally to the mesh deformation given a parameter value age=0.5age{=}0.5, as illustrated in Fig. 1. This use of interpolation constraints the shape space structure and helps producing topologically consistent shape deformations. The shape-adjusted model features a skeleton, composed of a set of bones connected along a kinematic tree. One bone is defined as root bone of the tree. To encode a pose for the model, the pose of the root bone is given, together with 3D rotations at each joint between connected bones, relative to their rest configuration. Given pose parameters, we apply forward kinematics to retrieve bone poses. We then deform the mesh using blend skinning, producing the final human mesh, shaped and posed accurately according to the input parameters. These steps are implemented using PyTorch [paszke2017automatic] and NVIDIA Warp [warp2022] to benefit from the automatic gradient backpropagation features of these libraries.
Self-collision detection.
To detect and prevent self-colliding poses, we test for intersecting faces belonging to different body parts using a bounding volume hierarchy for computational efficiency.
Interoperability.
To ensure interoperability with previous work, we define mappings between Anny and some existing mesh body models. These mappings serve mainly two purposes. The first is to empirically evaluate models trained with Anny on existing benchmarks; for that purpose we define a mapping to regress SMPL-X vertices from Anny meshes and vice versa. The second one is to generate synthetic 3D scenes with Anny annotations. Our synthetic data generation pipeline is built on Humgen3D [humgen3d], thus we also learn regressors for this body model. More specifically, we optimize sparse linear regressors $\bm{R}\in\mathbb{R}^{M\times N}$ to map vertex coordinates $\bm{v}_{j}\in(\mathbb{R}^{3})^{N}$ from a first body model to coordinates $\hat{\bm{v}}_{i}=\sum_{j}\bm{R}_{j,i}\bm{v}_{j}\in(\mathbb{R}^{3})^{M}$ of a different body model with a different topology. This is achieved by first fitting Anny to a set of meshes with the target connectivity (e.g. SMPL-X), and by initializing the regression coefficients $\bm{R}$ as the barycentric coordinates of the projection of target mesh vertices onto the source mesh. Coefficients of $\bm{R}$ are then refined jointly together with Anny parameters to minimize the mesh-to-mesh distance, while enforcing the left/right symmetry of the mapping. This optimization leads to a mean cyclic error of 3.23.2mm between Anny and SMPL-X (1.7mm resp. between Anny and Humgen3D).
4 Shape statistical modeling
Having a model of the distribution of human morphologies is useful for applications such as data synthesis, or as an optimization prior. Existing body models [loper2015smpl, park2025atlas] typically feature a latent space that encodes the shape distribution from their training set, e.g. using PCA for SMPL. These training sets consist of proprietary collections of 3D scans which required significant acquisition effort, yet they remain insufficient to represent the global population. Datasets such as CAESAR [robinette2002caesar] and SizeUSA [sizeusa] consist of scans of adults from western countries, and as such are not representative of morphological variations among all ages and all global population. In contrast, Anny can model a large diversity of morphologies thanks to its various phenotype parameters. This diversity covers common body shapes but also uncommon ones, such as 2.4 meters high individuals. We therefore calibrate the distribution of Anny parameters to obtain a shape distribution more resembling of the global population. We empirically define a bijective mapping between the age parameter of Anny and some morphological age in years. We then model distributions for major phenotype parameters as Beta distributions, conditioned on age and gender. These distributions are jointly calibrated to match the mean and standard deviation of reference growth standards and body mass index from the World Health Organization (WHO) [who2006growth], as shown in Figure 4.
5 Assessing 3D modeling capabilities
To assess how Anny can model the geometry of actual human meshes, we register it to scans of 3DBodyTex [saint20183dbodytex]. 3DBodyTex [saint20183dbodytex] is a dataset consisting of 400 3D scans of adults depicting 100 males and 100 female individuals in minimal clothing. We report qualitative results in Fig. 5. Anny approximates 3D points of the scans with an average point to mesh error of 2.4mm, excluding head and hands of the numerical evaluation due to the low scan quality in these regions, following an existing protocol [osman2022supr, park2025atlas]. Smaller fitting errors have been reported with other human mesh models, notably a 1.8mm mean vertex to vertex error using ATLAS [park2025atlas], however these models are specifically optimized to capture the geometry of individuals wearing minimal clothes. Anny models neither clothes nor the shape deformation they cause, yet is able to model these 3D scans with an accuracy sufficient for many applications. To assess modeling capabilities with children, we register Anny to three commercial scans of children from RenderPeople, and achieve a 2.7mm point to mesh error, excluding regions corresponding to the hairnet worn by the models of the numerical evaluation.
3D scan      Anny    Registration error
Mean error
Figure 5: 3D registration of Anny to scans of adults (top rows) and children (bottom).
6 Recovering Anny Human Meshes
Human Mesh Recovery (HMR) is a natural downstream task for assessing the representational power of a body model [multi-hmr2024, patel2025camerahmr]. Recent works [bedlam, tesch2025bedlam2] have shown that large-scale synthetic data can be as effective as, or even superior to, real data—whose annotations often contain inherent noise—for training HMR models. Having a body model interoperable with other meshes that can accommodate clothing enables to easily generate large-scale, diverse synthetic training data. In this paper, we leverage these properties and introduce Anny-One, a synthetic dataset of images with corresponding Anny annotations (Section 6.1). We then use Anny-One to train HMR models, that we evaluate on standard benchmarks (Section 6.2).
6.1 The Anny-One dataset
Our synthetic dataset, Anny-One, contains 800k images designed for training and evaluating Human Mesh Recovery models. It features realistic human meshes with diverse body shapes, poses, and appearances, situated in rich and varied scene contexts.
Figure 6: Random samples from the Anny-One dataset.
For each individual in the dataset, we generate both a ground-truth annotation using Anny and a corresponding HumGen3D character. Each character is randomly augmented with clothing and accessories sampled from the HumGen3D library, ensuring high visual diversity. These characters are then positioned within detailed, procedurally generated indoor scenes created using Infinigen Indoors [infinigen2024indoors]. Scene placement ensures no interpenetration between humans and surrounding objects. On average, each scene includes approximately five people, and we render up to 40 camera views per scene, with camera placement biased toward human-centric framing. We further enhance viewpoint diversity by rendering egocentric perspectives, upper-body close-ups, and hand-focused views for a subset of the data. The camera’s field of view is uniformly sampled between 30° and 130° to capture a wide range of spatial compositions.
Body poses are randomly sampled from AMASS [mahmood2019amass], while hand poses are independently drawn from GRAB[taheri2020grab]. Body shapes are sampled from a statistical distribution of human phenotypes derived from WHO population data (Section 5). To ensure physical plausibility, we apply a self-collision check to eliminate invalid or unrealistic configurations.
All images are rendered in Blender at a resolution of 1280×1280 pixels. Figure 6 shows representative samples from Anny-One, illustrating its diversity in body shapes, clothing, poses, and scene composition.
Overall, Anny-One provides a large-scale, visually rich, and statistically diverse resource for training HMR models in indoor environments.
6.2 HMR models
To empirically evaluate our Anny body model and Anny-One dataset, we rely on two recent state-of-the-art HMR models: HMR2.0 [goel2023humans] for the single-person setting and Multi-HMR [multi-hmr2024] for the multi-person setting. Both models are based on Vision transformers (ViTs). In both cases training code is available online, and we make minimal adaptations to predict Anny parameters instead of their original body models.
HMR2.0
[goel2023humans] relies on the ‘Huge’ variant of ViT with $16\times 16$ patch size, followed by a transformer decoder that takes as input a single learned token that cross-attends to all image tokens to output body model parameters. It takes human-centric image crops as inputs.
Multi-HMR
[multi-hmr2024] is also built on a ViT backbone. It is pre-trained with DINOv2 [dinov2] and released in various sizes (‘Small’, ‘Base’ and ‘Large’) with $14\times 14$ patches. The ViT backbone is also followed by a cross-attention-based decoder that processes backbone output tokens corresponding to detected people. The model takes full, uncropped images as input, and is trained to detect humans, regress expressive human meshes, and place them in the scene in 3D. We employ ViT-B ($448\times 448$) for ablation studies and ViT-L ($672\times 672$) for fair comparison with state-of-the-art methods.
Evaluation benchmarks.
We evaluate 3D mesh prediction accuracy on standard benchmarks: 3DPW [vonMarcard2018], EMDB [kaufmann2023emdb], Hi4D [yin2023hi4d], CMU-Toddler [cmu_panoptic_studio] and EHF [pavlakos2019smplx]. We report the commonly used Mean Per Joint Position Error (MPJPE) and Per Vertex Error (PVE), along with their Procrustes-aligned variants, following prior work [conman, multi-hmr2024]. We also report the Pair-PA-MPJPE, which measures the mean joint position error after Procrustes alignment of each pair of interacting humans, thereby evaluating the accuracy of their relative 3D poses. Although the CMU-Toddler dataset contains both adults and children, its limited diversity among children prevents extensive ablation studies. To better assess performance across diverse populations, we additionally use AGORA [agora], which includes both adults and children. While AGORA is synthetic, it remains the only dataset currently available for this purpose. Because the official validation set of AGORA does not contain any children, we re-define the train and val set of AGORA such that images with human scans from 3DPEOPLE constitute the new validation set composed of 2k images, ensuring an equal percentage of children in both training and validation splits.
6.3 HMR Results
To isolate the effect of our body model from that of our synthetic dataset, we first evaluate Anny by re-training existing HMR methods on established datasets. We then conduct experiments by training on Anny-One, optionally fine-tuning on the training set of the respective benchmarks. Finally we train using both Anny-One and standards HMR training data [kolotouros2019learning] including BEDLAM [bedlam], MS-COCO [lin2014microsoft], and MPII [andriluka14cvpr] to compare against state-of-the-art methods.
Model 3DPW EHF
Network Body MPJPE PA-MPJPE PVE PA-PVE
HMR2.0 [goel2023humans] SMPL-X 86.0 52.0 76.4 66.9
Anny 86.5 49.4 65.5 49.7
Multi-HMR [multi-hmr2024] SMPL-X 87.1 56.3 66.2 52.9
Anny 87.0 54.3 68.6 52.6
Table 1: Comparison of SMPL-X and Anny on 3DPW and EHF using HMR2.0 [goel2023humans] and Multi-HMR [multi-hmr2024] methods, trained on BEDLAM [bedlam] only.
Training AGORA-All AGORA-Kids
Pre-train Fine-tune Body PVE PA-PVE PVE PA-PVE
BEDLAM SMPL-X 140.8 71.3 186.0 64.2
Anny 136.7 71.3 175.6 63.6
Anny-One Anny 118.5 63.5 113.9 56.4
SMPL-X 89.7 60.8 99.5 50.6
SMIL-X 87.9 60.4 80.8 50.5
Anny 85.7 57.8 79.0 49.3
BEDLAM SMPL-X 78.2 50.3 96.5 45.6
BEDLAM SMIL-X 76.6 50.0 77.6 43.6
Anny-One Anny 72.8 48.2 69.3 41.5
Table 2: Effect of Model and Training Data on AGORA [agora] validation set (with diverse shapes from adults and children) using Multi-HMR [multi-hmr2024]. We use either BEDLAM or Anny-One for pre-training and SMPL-X/SMIL-X/Anny as body model.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Qualitative examples on real-world images sourced from Pexels [pexels] that contain both adults and children. We observe that our model outputs childlike body shapes for children, in terms of height and build.
Model 3DPW EMDB Hi4D CMU-Toddler
PA-MPJPE MPJPE PVE PA-MPJPE MPJPE PVE PA-MPJPE MPJPE Pair-PA-MPJPE MPJPE Pair-PA-MPJPE
AiOS [sun2024aios] 45.0 68.8 90.9 63.3 90.6 108.1 49.9 71.4 234.2 162.4 723.1
SAT-HMR [su2025sat] 52.7 81.0 94.5 71.0 112.9 126.7 61.2 88.2 85.5 153.9 654.8
Multi-HMR [multi-hmr2024] 46.9 69.5 88.8 48.5 73.7 87.1 49.8 67.8 80.6 153.6 638.9
Ours 41.8 71.5 83.2 48.5 71.5 83.4 48.7 66.6 80.0 102.1 263.8
Table 3: Comparison to state-of-the-art image-based methods. Quantitative results on multiple datasets comparing our approach against existing multi-person image-based methods. Ours corresponds to Multi-HMR with Anny. Lower values indicate better performance ($\downarrow$).
Anny vs SMPL-X: a simple body model is enough.
In order to compare Anny with SMPL-X [pavlakos2019smplx] in a model-agnostic manner, we adapt the implementations of two foundation models (HMR2.0 [goel2023humans] and Multi-HMR [multi-hmr2024]) to predict full-body meshes using either the Anny or SMPL-X body model. We train the HMR models on BEDLAM and evaluate them on the 3DPW and EHF test sets. To limit the computational cost of experiments, HMR2.0 is finetuned with a frozen backbone and without data augmentation.
Results are shown in Table 1. Across all metrics, Anny achieves comparable or superior performance to SMPL-X, despite its simpler and non data-driven design. This demonstrates that a compact, interpretable, and geometry-consistent body model such as Anny is sufficient for HMR and can serve as a drop-in replacement for SMPL-X.
Anny and Anny-One: modeling diverse body shapes.
We evaluate performance on AGORA, which is the only existing standard benchmark containing children, and report results in Table 2. First we train models on BEDLAM with SMPL-X or Anny. We see in the first and second rows of Table 2 that using the Anny head marginally improves performance, which may be due to the fact that Anny handles children more gracefully. Our proposed dataset consists of 800800k synthetic images, designed to contain humans with diverse body shapes, diverse backgrounds and changes in camera intrinsics. Training on Anny-One instead of BEDLAM (third row) brings substantial gains, in particular when evaluated on children.
We then consider training directly on AGORA, and compare Anny to both SMIL-X and SMPL-X. Both SMIL-X and Anny perform significantly better than SMPL-X, with a moderate advantage for Anny. This is consistent with the fact that they were both designed to handle children. Finally, we consider the full-data regime with pretraining on either BEDLAM or Anny-One, followed by finetuning on AGORA (last three rows of the table). We observe that in this large-scale regime, the combination of Anny-One and Anny significantly outperforms existing datasets and body models on AGORA.
Comparison to state-of-the-art methods.
We evaluate Multi-HMR trained with Anny on a large-scale mixture of datasets, including Anny and standard HMR training data (BEDLAM, MS-COCO, MPII), and compare against state-of-the-art multi-person HMR approaches. Results on 3DPW, EMDB, Hi4D, and CMU-Toddler (Table 3) show that our model achieves state-of-the-art performance across all benchmarks, consistently outperforming prior methods despite using a simpler, non data-driven, and interpretable body representation, which is of high interest for many downstream applications.
These results demonstrate that Anny scales effectively to large and diverse data, enabling robust reconstruction of adults and children alike across both in-the-wild and controlled multi-person scenarios. Anny achieves this performance without relying on data-driven models demonstrating that a simple, unified and interpretable human body model can scale effectively to modern HMR tasks.
Qualitative results.
We provide some qualitative examples of results of our approach on real images that contain both adults and children in Figure 7, showing high-quality human mesh recovery results with a single model thanks to Anny’s capacity of modeling a wide variety of body shapes.
7 Conclusion
We present Anny, a unified, differentiable, interpretable, and non data-driven human body model that continuously encode human shape variation across age, gender, and body type. Built entirely on open and procedural knowledge, Anny bridges the gap between artistic and statistical modeling, offering a transparent alternative to abstract latent spaces. Despite not being learned from 3D scans, Anny achieves competitive performance in fitting real scans, showing that accurate geometric modeling can be obtained purely from procedural priors and anthropometric calibration. Using Anny, we generate Anny-One, a large-scale dataset of 800k synthetic images with rich diversity in pose, shape, and scene context. Experiments show that HMR models trained with Anny achieve state-of-the-art performance across diverse benchmarks. By providing a free, interpretable and unified 3D model that covers the full human lifespan, we hope Anny will broaden the ecosystem of human-aware computer vision.